A 1200-kg car going 30 m/s applies its brakes and skids to rest. If the friction force between the sliding tires and the pavement is 6000 N, how far does the car skid before coming to rest ?


1
Expert's answer
2022-03-28T11:02:15-0400

The retardation of a car

"a=\frac{F}{m}=\frac{6000}{1200}=5\:\rm m/s^2"

The distance traveled by a car to the rest

"d=\frac{v_0^2}{2a}=\frac{30^2}{2*5}=90\:\rm m"